toyo Manufacturing has 1,000,000 shares of common stock outstanding at a market price of $40.20 a share. Last month, the company paid an annual dividend in the amount of $2.34 per share. The dividend growth rate is 3%. Toyo Manufacturing also has 25,000 bonds outstanding with a face value of $1,000 per bond. The bonds carry a 6% coupon, pay interest annually, and mature in 10 years. The bonds are selling at 98% of face value. The company's tax rate is 34%. What is Toyo Manufacturing's weighted average cost of capital?
      
A6.87%      
B6.45%      
C7.16%      
D7.23%      
E7.54%
Answer #1
As per DDM
Price = recent dividend* (1 + growth rate )/(cost of equity - growth rate)
40.2 = 2.34 * (1+0.03) / (Cost of equity - 0.03)
Cost of equity% = 9
MV of equity=Price of equity*number of shares outstanding
MV of equity=40.2*1000000
=40200000
MV of Bond=Par value*bonds outstanding*%age of par
MV of Bond=1000*25000*0.98
=24500000
MV of firm = MV of Equity + MV of Bond
=40200000+24500000
=64700000
Cost of debt
                  K = N
Bond Price =∑ [(Annual Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =10
980 =∑ [(6*1000/100)/(1 + YTM/100)^k]     +   1000/(1 + YTM/100)^10
                   k=1
YTM = 6.275291107
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 6.275291107*(1-0.34)
= 4.14169213062
Weight of equity = MV of Equity/MV of firm
Weight of equity = 40200000/64700000
W(E)=0.6213
Weight of debt = MV of Bond/MV of firm
Weight of debt = 24500000/64700000
W(D)=0.3787
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=4.14*0.3787+9*0.6213
WACC% = 7.16